# Preference Tuning with Human Feedback on Language, Speech, and Vision Tasks: A Survey

### v1

@@ Abstract
Preference tuning is a crucial process for aligning deep generative models with human preferences. This survey offers a thorough overview of recent advancements in preference tuning and the integration of human feedback. The paper is organized into three main sections: 1) introduction and preliminaries: an introduction to reinforcement learning frameworks, preference tuning tasks, models, and datasets across various modalities: language, speech, and vision, as well as different policy approaches, 2) in-depth examination of each preference tuning approach: a detailed analysis of the methods used in preference tuning, and 3) applications, discussion, and future directions: an exploration of the applications of preference tuning in downstream tasks, including evaluation methods for different modalities, and an outlook on future research directions. Our objective is to present the latest methodologies in preference tuning and model alignment, enhancing the understanding of this field for researchers and practitioners. We hope to encourage further engagement and innovation in this area.
@@ 1 Introduction
Learning from human feedback is a crucial step in aligning generative models with human preferences to generate output that closely resembles human speech and writing. Despite the powerful learning capabilities of generative models in self-supervised learning, these models frequently misinterpret instructions, leading to hallucinations in generation (Ji et al., 2023a; Yao et al., 2023a). Additionally, ensuring the safety of the generated content remains a significant challenge for these models. Extensive research on preference tuning using human feedback has demonstrated that adversarial samples can be utilized to jailbreak systems (Rando and Tramèr, 2023; Wei et al., 2024). Ideally, generative models need to be controlled to ensure that their outputs are safe and do not cause harm. Models often exhibit unintended behaviors, such as fabricating facts (Chen and Shu, 2023; Sun et al., 2024), producing biased or toxic text (Hartvigsen et al., 2022), or failing to follow user instructions (Ji et al., 2023b; Tonmoy et al., 2024). Additionally, maintaining the privacy of data is crucial to ensure the safe operation of models and protect user privacy (Brown et al., 2022). In the text-to-image generation task, large-scale models often struggle to produce images that are well-aligned with text prompts (Feng et al., 2022), particularly in compositional image generation (Liu et al., 2022; Lee et al., 2023) and coherent generation (Liu et al., 2023a). Similarly, in text-to-speech tasks, Zhang et al. (2024a); Chen et al. (2024a) integrate subjective human evaluation into the training loop to better align synthetic speech with human preferences.
@@ 2.2 Taxonomy
            (Dang et al., 2024) $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$
@@ 3 Preference Tuning
Figure 4: Preference Tuning methods for online algorithms, such as RLHF (PPO/REINFORCE), Online DPO, and SFT-like, and offline methods, such as DPO.
@@ 3.1.2 Reward Modeling
GRM-llama3-8B-sftreg (Yang et al., 2024) 8B Llama3 Preference 700K
GRM-llama3-8B-distill (Yang et al., 2024) 8B Llama3 Preference 700K
@@ 3.1.4 Joint Training
WildChat (Zhao et al., 2024) 652k $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@ 3.2.1 SFT Datasets
Taori et al. (2023) propose Alpaca, a dataset with demonstrations generated using OpenAI’s GPT-3 text-davinci-003 model. The instruction data can be used to conduct instruction tuning for LLMs and allow them to follow instruction better. A version of Alpaca dataset with Chain-of-Thought (CoT) (Wei et al., 2022) and it is introduced to further improve the LLM’s reasoning ability. Multi-turn datasets generated using LLMs are also created, such as ChatAlpaca, UltraChat (Ding et al., 2023), and WildChat (Zhao et al., 2024).
@@ 3.2.2 Human Preference Alignment Datasets
mBART (Liu et al., 2020) 406M Multi. (25), Multi. (50) Enc-Dec $\checkmark$ $\times$
@@ 5.1.2 Rejection Sampling Optimization (RSO)
for any convex function $f$, like in Table 12. GPO further provides an analysis of this formulation from an policy improvement and policy regularization tradeoff. Applying Taylor Expansion of the form above yields:
@@ 5.2.1 Diffusion-DPO
Diffusion-DPO (Wallace et al., 2024) is adapting DPO to diffusion models where it uses a fixed dataset where each example contains a prompt and a pairs of images generated from a reference model with human label. Similar to RL for diffusion, the goal is still to align the base diffusion models to human preferences. The derivation is similar to RL framework for diffusion in DDPO and DPOK, and also DPO for Language Models:
@@ 6 Combined Policies and Sampling-Agnostic Alignment
In this section, we explore a different approach to human preference tuning by integrating off-policy and on-policy methods. This combined approach has proven to be highly effective and complementary in enhancing the alignment quality of the model. Additionally, we introduce sampling-agnostic alignment methods that can be applied to both off-policy and on-policy approaches.
@@ 6.1 ExPO
ExPO (Zheng et al., 2024a) provides a simple and training-free method for enhancing the alignment of large language models (LLMs) with human preferences. The core insight behind ExPO is that a model trained with DPO/RLHF can be viewed as an interpolation between two models with differing strengths. By leveraging this concept, one can potentially extrapolate a stronger model if the other two models are available. Specifically, if we denote the model $\pi_{\text{ExPO}}$ as the interpolation of two other models, $\pi_{a}$ and $\pi_{b}$, which may be trained using different alignment methods and datasets, ExPO assumes that combining these models will yield improved alignment. The stronger, better-aligned model $\pi_{ExPO}$ is then obtained by extrapolating from the weights of these two relatively weaker models, as formulated below:
@@ 6.3 Reinforced Token Optimization (RTO)
RTO (Zhong et al., 2024) derives a token-wise reward function from preference data and conducts policy optimization using this learned reward signal. Broadly, RTO involves two primary steps: (1) learning a token-wise reward from preference data, and (2) optimizing this reward through RL training methods like PPO. Theoretically, the algorithm is framed as an MDP problem.
Theoretical Version
It focuses on the offline setting assuming it has an offline dataset $\mathcal{D}=\{(\tau^{w},\tau^{l})\}$ that contains several trajectory pairs, where $\tau^{w}=\{(s_{h}^{w},a_{h}^{w})\}_{h=1}^{H}$ is preferred over $\tau^{l}=\{(s_{h}^{l},a_{h}^{l})\}_{h=1}^{H}$. Each pair of trajectories shares the same initial state (i.e., $s_{1}^{w}=s_{1}^{l}$), but differs in the subsequent tokens. They assume that the reward function is linear, and the results are extended to general function approximation. RTO computes the $\theta_{mle}$ based on $\mathcal{D}$ by maximizing the log likelihood and calculates the pessimistic reward $\hat{r}$ via token-wise reward learning. The output of the algorithm is policy $\hat{\pi}$.
@@
The key challenge in implementing RTO lies in learning the token-wise reward to be optimized from the offline data. In the most popular frameworks outlined in InstructGPT (Ouyang et al., 2022), Claude (Bai et al., 2022a), and LLaMA2 (Touvron et al., 2023a) projects replace the last layer of the LLM with a linear layer for a scalar output and maximize the log-likelihood. However, this approach gives only a sentence-level reward. To bridge the gap in the literature, they present the practical version of RTO, which features a novel calculation of token-wise reward. Their key observation is that, given a trajectory $\tau=\{(s_{h},a_{h})\}_{h=1}^{H}$, RTO is defined the following:
@@
where the first equality uses the closed-form of optimal policy $\pi_{\beta_{\text{reg}}}^{*}(a|s)=\exp\{(Q_{\beta_{\text{reg}}}^{*}(s,a)-V_{\beta_{\text{reg}}}^{*}(s))/\beta_{\text{reg}}\}$, and the second equality follows from the fact that $Q_{\beta_{\text{reg}}}^{\pi}(s,a)=r_{\beta_{\text{reg}}}(s,a)+\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[V_{\beta_{\text{reg}^{\pi}}}(s^{\prime})]$ with $r_{\beta_{\text{reg}}}(s,a)=r(s,a)+\beta_{\text{reg}}\log\pi_{\mathrm{ref}}(a|s)$. RTO focuses on the typical LLM generation scenario where the transition kernel is deterministic. Then, $(\star)=0$, yielding that
@@
as the token-wise reward used by RTO, where $\beta_{\text{reg}}^{1}$ and $\beta_{\text{reg}}^{2}$ are tuning parameters, and $\pi$ is the current policy to be updated. In the last step, RTO uses $\pi_{\mathrm{dpo}}$, the policy learned by DPO, as a proxy for the unknown $\pi_{\beta_{\text{reg}}}^{*}$. Finally, RTO employs PPO to optimize the token-wise reward $r_{\mathrm{rto}}$.
@@ 7.1.1 AlpacaEval
AlpacaEval (Dubois et al., 2024) win rate (against GPT4) is an LLM-based automatic evaluation that has high-level agreement to human. To further improve the fairness of the evaluation and address the verbosity of issue of GPT4 as a judge, Dubois et al. (2024) introduce a length-controlled version of AlpacaEval that aims to conduct measurement with outputs with similar lengths. The metric is used in AlpacaEval calculates win-rates for models across a variety of NLP tasks to measure of model capabilities compared to a baseline by using an LLM judge.
AlpacaEval 2.0
The judge uses GPT4-Turbo to replace GPT-3 based model “text-davinci-003” in the 1.0 version, which makes it more challenging and have a metric that better reflects the current SOTA model.

### v2

@@ Abstract
Preference tuning is a crucial process for aligning deep generative models with human preferences. This survey offers a thorough overview of recent advancements in preference tuning and the integration of human feedback. The paper is organized into three main sections: 1) introduction and preliminaries: an introduction to reinforcement learning frameworks, preference tuning tasks, models, and datasets across various modalities: language, speech, and vision, as well as different policy approaches, 2) in-depth exploration of each preference tuning approach: a detailed analysis of the methods used in preference tuning, and 3) applications, discussion, and future directions: an exploration of the applications of preference tuning in downstream tasks, including evaluation methods for different modalities, and an outlook on future research directions. Our objective is to present the latest methodologies in preference tuning and model alignment, enhancing the understanding of this field for researchers and practitioners. We hope to encourage further engagement and innovation in this area.
@@ 1 Introduction
Learning from human feedback is a crucial step in aligning generative models with human preferences to generate output that closely resembles human speech and writing. Despite the powerful learning capabilities of generative models in self-supervised learning, these models frequently misinterpret instructions, leading to hallucinations in generation (Ji et al., 2023a; Yao et al., 2023a). Additionally, ensuring the safety of the generated content remains a significant challenge for these models. Extensive research on preference tuning using human feedback has demonstrated that adversarial samples can be utilized to jailbreak systems (Rando and Tramèr, 2023; Wei et al., 2024). Ideally, generative models need to be controlled to ensure that their outputs are safe and do not cause harm. Models often exhibit unintended behaviors, such as fabricating facts (Chen and Shu, 2023; Sun et al., 2024), producing biased or toxic text (Hartvigsen et al., 2022), or failing to follow user instructions (Ji et al., 2023b; Tonmoy et al., 2024). Additionally, maintaining the privacy of data is crucial to ensure the safe operation of models and protect user privacy (Brown et al., 2022). In the text-to-image generation task, large-scale models often struggle to produce images that are well-aligned with text prompts (Feng et al., 2022), particularly in compositional image generation (Liu et al., 2022; Lee et al., 2023), object recognition (Qiao et al., 2024), and coherent generation (Liu et al., 2023a). Similarly, in text-to-speech tasks, Zhang et al. (2024a); Chen et al. (2024a) integrate subjective human evaluation into the training loop to better align synthetic speech with human preferences.
@@ 2.2 Taxonomy
            Multi-Ling RLHF (Dang et al., 2024) $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$
@@
            GRPO (Shao et al., 2024) $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
@@
            PRDP (Deng et al., 2024) $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
@@
      VPO (Chen et al., 2024b) $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
@@
      RainbowPO (Zhao et al., 2024a) $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
@@ 3 Preference Tuning
Figure 4: Preference Tuning methods for online algorithms, such as RLHF, Online DPO, and SFT-like, and offline methods, such as DPO.
@@ 3.1.2 Reward Modeling
GRM-llama3-8B-sftreg (Yang et al., 2024a) 8B Llama3 Preference 700K
GRM-llama3-8B-distill (Yang et al., 2024a) 8B Llama3 Preference 700K
@@
Multi-Model
MetaMetrics-RM (Winata et al., 2024a) Multiple Multiple Skywork Preference Data and AllenAI Preference Data
@@
Multi-Model Reward Model
MetaMetrics (Winata et al., 2024a) is a method to combine multiple existing reward models into a more powerful reward model by calibrating them using the preference data. The method is a systematic way to identify reward models that can be used complementary without blindly use the models. There are two methods introduced to calibrate the models using Bayesian optimization and boosting method. Thus, the approach is highly efficient and they are aspect-agnostic, thus allowing flexibility to use them in any preference data.
@@ 3.1.4 Joint Training
M2Lingual (Maheshwary et al., 2024) 174k $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
@@
WildChat (Zhao et al., 2024b) 652k $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@ 3.2.1 SFT Datasets
Taori et al. (2023) propose Alpaca, a dataset with demonstrations generated using OpenAI’s GPT-3 text-davinci-003 model. The instruction data can be used to conduct instruction tuning for LLMs and allow them to follow instruction better. A version of Alpaca dataset with Chain-of-Thought (CoT) (Wei et al., 2022) and it is introduced to further improve the LLM’s reasoning ability. Multi-turn datasets generated using LLMs are also created, such as ChatAlpaca, UltraChat (Ding et al., 2023), and WildChat (Zhao et al., 2024b).
@@ 3.2.2 Human Preference Alignment Datasets
mBART (Liu, 2020) 406M Multi. (25), Multi. (50) Enc-Dec $\checkmark$ $\times$
@@ 4.5.4 AlignProp
4.5.5 Proximal Reward Difference Prediction
PRDP (Deng et al., 2024) proposed a loss for matching likelihood difference with reward difference for fine-tuning diffusion models, inspired by DPO. Notice that, (same as derivation in DPO), for any two generations $x^{1}_{0}$ and $x^{2}_{0}$, the optimal policy (KL-regularized reward) yields:
$\log\frac{\pi_{\theta^{\star}}\left(x^{1}_{0}\mid\mathbf{c}\right)}{\pi_{\text{ref }}\left(x^{1}_{0}\mid\mathbf{c}\right)}-\log\frac{\pi_{\theta^{\star}}\left(x^{2}_{0}\mid\mathbf{c}\right)}{\pi_{\text{ref }}\left(x^{2}_{0}\mid\mathbf{c}\right)}=\frac{r\left(x_{0}^{1},\mathbf{c}\right)-r\left(x_{0}^{2},\mathbf{c}\right)}{\beta_{\text{reg}}}$ (31)
thus PRDP proposes to minimize the MSE error between LHS with $\theta$ (replacing $\theta^{*}$) and RHS. The objective is $\mathcal{L}_{\mathrm{PRDP}}\left(\pi_{\theta};\pi_{\mathrm{ref}}\right):=$
$\mathbb{E}_{c\sim\mathcal{D},\left(x^{1},x^{2}\right)\sim\pi_{\theta}(\cdot\mid c)}\left(\beta_{\text{reg}}\log\frac{\pi_{\theta}\left(x_{0}^{1}\mid x\right)}{\pi_{\text{ref }}\left(x_{0}^{1}\mid x\right)}-\beta_{\text{reg}}\log\frac{\pi_{\theta}\left(x_{0}^{2}\mid x\right)}{\pi_{\text{ref }}\left(x_{0}^{2}\mid x\right)}-(r(x^{1}_{0})-r(x^{2}_{0}))\right)^{2},$ (32)
Furthermore, they also employ proximal updates (clipping the ratios and optimizing a proximal objective) for stable training of (32), in the same spirit of PPO.
Similar works include Yang et al. (2024b), which applies the idea of dense reward to DPO-style explicit-reward-free approach on text-to-image diffusion models, so as to suit better to diffusion models’ generation hierarchy.
@@ 5 Offline Alignment
RainbowPO $-\log\sigma\left(\phi(x)\left[\frac{\beta_{\text{reg}}}{\left|y^{w}\right|}\log\frac{\pi_{\theta}\left(y_{w}\mid x\right)}{\pi_{\alpha}\left(y_{w}\mid x\right)}-\frac{\beta_{\text{reg}}}{\left|y^{l}\right|}\log\frac{\pi_{\theta}\left(y_{l}\mid x\right)}{\pi_{\alpha}\left(y_{l}\mid x\right)}\right]\right)$
@@ 5.1.2 Rejection Sampling Optimization (RSO)
for any convex function $f$, like in Table 12. GPO further provides an analysis of this formulation from an policy improvement and policy regularization trade-off. Applying Taylor Expansion of the form above yields:
@@ 5.1.10 SimPO
5.1.11 RainbowPO
Inspired by the paper Rainbow on improving DQN for better performance, RainbowPO (Zhao et al., 2024a) demystifies the effectiveness of existing DPO variants by categorizing their key components into several broad directions, and integrate the identified effective components into a single cohesive objective:
$\mathcal{L}_{\text{RainbowPO}}\left(\pi_{\theta};\pi_{\text{ref }}\right)=-\underset{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}}{\mathbb{E}}f\left[\phi(x)\left(\frac{\beta}{\left|y^{w}\right|^{\eta}}\log\frac{\pi_{\theta}\left(y_{w}\mid x\right)}{\pi_{\alpha}\left(y_{w}\mid x\right)}-\frac{\beta}{\left|y^{l}\right|^{\eta}}\log\frac{\pi_{\theta}\left(y_{l}\mid x\right)}{\pi_{\alpha}\left(y_{l}\mid x\right)}\right)\right],$ (60)
in which $\eta\in\{0,1\}$, and $\pi_{\alpha}$ is referred to a mixing policy mechanism they propose for formulating a better reference policy by mixing policy $\pi_{\text{ref}}$ and $\pi_{\gamma}$, defined as:
$\pi_{\alpha}(y\mid x)\propto\pi_{\mathrm{ref}}^{\alpha}(y\mid x)\cdot\pi_{\gamma}^{1-\alpha}(y\mid x),$ (61)
and $\pi_{\gamma}$ is a policy which assumes to exist (which can be understood as the reference policy taken by SimPO (Meng et al., 2024)), such that the model is perfect at distinguishing the preference pairs in the dataset:
$\pi_{\gamma}\left(y_{w}\mid x\right)^{1/\left|y^{\omega}\right|}/\pi_{\gamma}\left(y_{l}\mid x\right)^{1/\left|y^{l}\right|}=\exp(\gamma),$ (62)
for any prompt $x$. Zhao et al. (2024a) show that optimizing such generic objective can yield the best performance on downstream task of tuning Llama3-8B-Instruct for instruction-following capabilities, benefiting from composition of effective elements.
@@ 5.2.1 Diffusion-DPO
Diffusion-DPO (Wallace et al., 2024) is adapting DPO to diffusion models. It uses a fixed dataset and each example contains a prompt and a pairs of images generated from a reference model with human label. Similar to RL for diffusion, the goal is still to align the base diffusion models to human preferences. The derivation is similar to RL framework for diffusion in DDPO and DPOK, and also DPO for Language Models:
@@ 6 Combined Policies and Sampling-Agnostic Alignment
In this section, we explore some other directions proposed in literature for improving the effectiveness of human preference tuning. We discuss ExPO (Zheng et al., 2024a), which proposed that combining two aligned models by extrapolating from their weights could enhance the alignment quality of the model; we discuss P3O (Fakoor et al., 2020), which utilized both on-policy and off-policy sampling; we also introduce sampling-agnostic alignment methods that can be applied to both off-policy and on-policy approaches.
@@ 6.1 ExPO
ExPO (Zheng et al., 2024a) provides a simple and training-free method for enhancing the alignment of large language models (LLMs) with human preferences. The core insight behind ExPO is that a model trained with DPO/RLHF can be viewed as an interpolation between two models with differing strengths. By leveraging this concept, one can potentially extrapolate a stronger model if the other two models are available. Specifically, if we denote the model $\pi_{\text{ExPO}}$ as the interpolation of two other models, $\pi_{a}$ and $\pi_{b}$, which may be trained using different alignment methods and datasets, ExPO assumes that combining these models will yield improved alignment. The stronger, better-aligned model $\pi_{ExPO}$ is then obtained by extrapolating from the weights of these two relatively weaker models (which is reminiscent to Model Soups (Wortsman et al., 2022)), as formulated below:
@@ 6.3 Reinforced Token Optimization (RTO)
Standard RLHF and DPO’s reward models are all based on the whole generation, thus the whole pipeline is in some sense closer to bandit instead of classical MDP based RL. Inspired by that nature of auto-regressive models is next token prediction , RTO (Zhong et al., 2024) derives a token-wise reward function from preference data and conducts policy optimization using this learned reward signal. Broadly, RTO formulates the optimization problem as an MDP and involves two primary steps: (1) learning a token-wise reward from preference data, and (2) optimizing this reward through RL training methods like PPO.
Theoretical Version.
Consider the offline setting by assuming that we have an offline dataset $\mathcal{D}=\{(\tau^{w},\tau^{l})\}$ that contains several trajectory pairs, where $\tau^{w}=\{(s_{h}^{w},a_{h}^{w})\}_{h=1}^{H}$ is preferred over $\tau^{l}=\{(s_{h}^{l},a_{h}^{l})\}_{h=1}^{H}$. Each pair of trajectories shares the same initial state (i.e., $s_{1}^{w}=s_{1}^{l}$), but differs in the subsequent tokens. RTO computes the maximum likelihood estimator $\theta_{mle}$ based on $\mathcal{D}$ by maximizing the log likelihood and calculates the pessimistic reward $\hat{r}$ via token-wise reward learning. The output of the algorithm is policy $\hat{\pi}$.
@@
Similar to learning the reward model in RLHF, the key challenge left is to learn the token-wise reward from the offline data. For sentence level reward, popular frameworks outlined in InstructGPT (Ouyang et al., 2022), Claude (Bai et al., 2022a), and LLaMA2 (Touvron et al., 2023a) replace the last layer of the LLM with a linear layer for a scalar output and maximize the log-likelihood, which thus cannot be naively used for token-level reward. RTO observes that, given a trajectory $\tau=\{(s_{h},a_{h})\}_{h=1}^{H}$, denoting $\pi_{\beta_{\text{reg}}}^{*}(a|s)=\exp\{(Q_{\beta_{\text{reg}}}^{*}(s,a)-V_{\beta_{\text{reg}}}^{*}(s))/\beta_{\text{reg}}\}$ as the optimal policy, the KL regularization can be rewritten as:
@@
in which the second equality follows from the fact that:
$Q_{\beta_{\text{reg}}}^{\pi}(s,a)=r_{\beta_{\text{reg}}}(s,a)+\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[V_{\beta_{\text{reg}^{\pi}}}(s^{\prime})],$ (76)
with $r_{\beta_{\text{reg}}}(s,a)=r(s,a)+\beta_{\text{reg}}\log\pi_{\mathrm{ref}}(a|s)$. RTO focuses on the typical LLM generation scenario where the transition kernel is deterministic. Then, $(\star)=0$, yielding that
@@
as the token-wise reward used by RTO, where $\beta_{\text{reg}}^{1}$ and $\beta_{\text{reg}}^{2}$ are tuning parameters, and $\pi$ is the current policy to be updated. In the last step, RTO uses $\pi_{\mathrm{dpo}}$, the policy learned by DPO, as a proxy for the unknown optimal $\pi_{\beta_{\text{reg}}}^{*}$. Then RTO employs PPO to optimize the model with respect to the token-wise reward $r_{\mathrm{rto}}$. The idea of transformation from sequence level preferences to token level guidance also appeared in an earlier work by Yang et al. (2024c).
@@ 7.1.1 AlpacaEval
AlpacaEval (Dubois et al., 2024) win rate (against GPT4) is an LLM-based automatic evaluation that has high-level agreement to human. To further improve the fairness of the evaluation and address the verbosity of issue of GPT4 as a judge, Dubois et al. (2024) introduce a length-controlled version of AlpacaEval that aims to conduct measurement with outputs with similar lengths. The metric is used in AlpacaEval calculates win-rates for models across a variety of NLP tasks to measure of model capabilities compared to a baseline by using an LLM judge. AlpacaEval 2.0: The judge uses GPT4-Turbo to replace GPT-3 based model “text-davinci-003” in the 1.0 version, which makes it more challenging and have a metric that better reflects the current SOTA model.